OpenStudy (anonymous):

Kathy is 6 years younger than Bill. 12 years ago, Bill was twice as old as Kathy was then. How old are they now? Explain how you got the equation!

OpenStudy (anonymous):

what would make you times by 6

OpenStudy (anonymous):

x + 6 = y So twelve years ago, 2x = y x+6 =2x -12 Solve for x (girls age) then you plug into y (boys age). x= 18. So therefore 18 + 6 = y

OpenStudy (anonymous):

k=b-6 12 years ago this was true: b=2k 12 years later, today, b+12=2k+12 Solve the following for k and b: {k = b - 6, b + 12 = 2 k + 12} {b = 12, k = 6}